OpenStudy (anonymous):

When Sean left for college, his parents decided to move from their two thousand four hundred square-foot home into a smaller house. If the size of their house decreased by forty percent, how many square feet is their new house? A)nine hundred sixty square feet B)one thousand four hundred forty square feet C)one thousand eight hundred square feet D)two thousand three hundred sixty square feet

OpenStudy (jdoe0001):

their original home was \(\bf 2,400 ft^2\) their new home is 40% smaller than that so their new home is 40% of 2,400 \(\bf \textit{n% of m is }\cfrac{n}{100}\cdot m\qquad thus\\ \quad \\ \textit{40% of 2,400 is }\quad \cfrac{40}{100}\cdot 2,400\\ \quad \\ \textit{their new smaller home is }2,400 - \left(\cfrac{40}{100}\cdot 2,400\right) \)

OpenStudy (jdoe0001):

so their new home is 2,400- (40% of 2,400) I meant